# JWST/NIRSpec insights into the circumnuclear region of Arp 220: A detailed kinematic study

### v1

@@ 4.2 Hourglass outflow from the W nucleus
As shown in Fig. 1, both the ionized and molecular gas exhibit a shell-shaped morphology at high red-shifted ( km s-1) velocities in the region north of the W nucleus. Additionally, the hot molecular gas shows a blue-shifted shell-like structure in the region south-west of the same nucleus. The blue- and red-shifted gas forms an ‘hourglass’ (HG)-like shape, compatible with an outflow launched from the W nucleus. This structure is even clearer in Fig. 5c, showing the sum of H2 spectral channel maps at km s-1 and km s-1 (with respect to the systemic velocity of the W nucleus), used to define our mask (white contours in the figure).
Within the HG region, we fitted simultaneously the hot molecular and the ionized gas using three Gaussian components, taking into account the large- and small-scale gas rotations as well as the outflow. To assign a Gaussian component to the HG outflow (HGO), we discarded the components with absolute velocities km s-1. Among the remaining ones, we attributed to the outflow the component with highest absolute velocity. The threshold (180 km s-1) represents a compromise: with a smaller velocity threshold, we would have included in the outflow also part of the rotational disc at larger scales; with a higher value, we would have missed part of the contribution of the outflow.
Figure 8 shows the flux, velocity and velocity dispersion of the gas component attributed to the outflow. The flux is dominated by the hot molecular gas over the ionised one. The velocity field is rather symmetrical with respect to the W nucleus and varies between and 400 km s-1. The highest observed velocity is located to the east in the redshifted shell, corresponding to an increase in molecular gas emission line flux. It is unclear whether this high-velocity gas is still part of the HGO, or whether it may represent a red-shifted counterpart of the conical SE outflow associated with the E nucleus (Sect. 4.1).
@@
4.3 North-West Bubble
We identified a large scale bubble in the north west of the NIRSpec FoV, mostly visible in the [Fe ii] transitions (Fig. 1). Figure 5d displays the sum of spectral channel maps at km s-1 (with respect to the systemic velocity of E nucleus) of Pa and H2. We created the mask for the NW bubble (NWB) by selecting those spaxels with enhanced flux in these velocity channels, excluding the region around the W nucleus (white curve in the figure). This nuclear region was already included in Sect.4.2, but since its outflow kinematics are slightly different from those of the HG outflow, this region will be discussed in Sect. 4.4.
The emission that contributes to the line profiles in this mask comes from the bubble and the large-scale rotating disc. Consequently, we fitted the emission lines with two Gaussian components. No specific boundaries on velocity and velocity dispersion were used during the fit. We assigned the most blue-shifted component to the outflow. This Gaussian component also exhibits a higher velocity dispersion with respect to the one associated with the disk.
@@ 4.3 North-West Bubble
Figure 11 shows the fit of Pa and H2 in three spaxels extracted from the bubble. The black dashed curve represents the spectra from the original cube, highlighting the component (red-shaded area) attributed to the HGO that we have already isolated and subtracted. The red curves represent the NWB component, while the orange curves correspond to the other components used in the fit. In all panels, the nature of the Gaussian components is clear: the outflow component is blue-shifted and broader than the disk component.
@@ 4.4 Compact outflow from the W nucleus
There is another kinematic feature in the W nuclear region that has not been isolated yet. It is a nuclear outflow (WNO) that is visible in the ionised gas phase (see also Perna et al. 2024). However, it affects just a few spaxels and cannot be isolated in the moment maps (Fig. 4) nor in the velocity channel maps (e.g. Fig. 1). Therefore, once all the features described before (i.e. the SEO, the HGO, and the NWB) have been subtracted, we fitted the residual cube with two components to model the WNO and the perturbed rotating gas (i.e. large-scale rotation over the entire FoV, or small-scale rotating nuclear discs).
@@
We attributed to the WNO the components with ¿ 150 km s-1 within a radius of 0.4” (150 pc) around the nucleus. In the more external regions, we associated both the components to the perturbed rotating gas. Figure 12 shows the three moments of the components attributed to the perturbed disc in H2 and Pa (top and central panels) and of the WNO (bottom panels). The WNO kinematics are characterized by blue-shifted velocities up to km s-1 and a velocity dispersion of km s-1.
@@
Figure 13 shows the spectrum extracted from a random spaxel (marked in the bottom panel of Fig. 12) to illustrate the fit decomposition for the ionised and hot molecular gas phases. The differences in the line profile of the two gas phases are evident: the broad, blue-shifted wing that is observed in Pa is not observed in the molecular gas.
4.5 Perturbed Rotating Disc
@@
The moment maps of hot molecular and ionised gas shown in Fig. 4 already suggested a large-scale rotation pattern, although strongly affected by several outflow features. These additional kinematic components were isolated and subtracted in Sects. 4.1, 4.2, 4.3 and 4.4. The fitting procedure used to separate the compact nuclear outflow (WNO) component from the rotating disk is described in the previous section (Sect. 4.4). After the subtraction of the components associated with the WNO, we assigned to the perturbed rotating gas all the components having km s-1. In fact, the components with higher velocity dispersion might represent residuals of outflowing gas that we did not model in the previous steps. Figure 12 (top and central panels) shows the moment maps of the isolated perturbed rotating gas, with contributions from both circum-nuclear discs in the vicinity of the E and W nuclei and the large-scale disc. The Moment-1 velocities are defined with respect to , which is more appropriate for the kpc-scale disc (see Sect. 3.3).
@@
We observe a symmetric velocity gradient with respect to the E nucleus, with velocities from km s-1 (north-east) to km s-1 (south-west). The Moment-1 map in the north-west part of the FoV is instead disturbed due to the non-perfect subtraction of HGO and NWB contributions. Moreover, significant deviations from a large-scale rotational patter are also observed in the vicinity of the W nucleus: this is due to the presence of a small-scale disc surrounding the W nucleus (see also e.g. Scoville et al. 2017). We calculated the position angle (PA)111The position angle is measured on the receding half of the galaxy, taken anticlockwise from the north direction on the sky (east of north). of the major kinematic axis of the large-scale disc by fitting the Moment-1 map of Pa and for H2 with the FIT KINEMATIC PA code (Cappellari et al., 2007; Krajnović et al., 2011), obtaining 517° and 397°, respectively. In Sect. 4.6 we will use the average of these two values to construct the PV diagram.
@@
The Moment-2 of the perturbed rotating disc reaches values up to 150 km s-1, significantly lower than in the moment 2 map obtained before the subtraction of the outflowing features (Fig. 4). Moreover, the kinematics of Pa are more perturbed compared to that of the hot molecular gas, as also seen in Armus et al. (2023).
@@
Figure 14 shows three single-spaxel spectra extracted from the HG region (B), the SEO cone (C), and an overlapping region of the HG and NW bubble (D) (See Fig. 12). After the subtraction of the outflow components, the line profiles are now rather symmetrical; the second component used in the fit procedure (red dashed curves) either is much fainter than the first component, or shows similar velocity shifts and line widths. Therefore, the attribution of both components to the disc is well justified.
@@ 4.6 Stellar Kinematics
Figure 15 presents the stellar kinematics maps obtained with pPXF. In particular, they are created from the analysis of the G235H cube, that is less affected by dust attenuation than the G140H cube and shows more prominent CO stellar absorption features in the wavelength range 2.2–2.4 m (Fig. 2). In Fig. 20 we show that the maps derived from the G235H and G140H gratings are consistent. As shown in the inset panel of Fig. 15, the stellar velocity field of Arp 220 presents two nuclear discs rotating in opposite directions separated by 1”, with a velocity offset of 120 km s-1 (see also Wheeler et al. 2020; Sakamoto et al. 1999). The stellar velocity dispersion is higher in the nuclear positions, and then decreases throughout a region crossing the two nuclei, in correspondence with the bright elongated region observed in the Pa Moment-0 map (Fig. 12). Moreover, in the E nucleus, the velocity dispersion increases towards the east, while in the W nucleus it increases in the south direction. In the outermost regions north of the two nuclei, the velocity dispersion reaches about 200 km s-1, consistent with the values found in Perna et al. (2020).
In the next two sub-sections we compare the stellar and gas kinematics making use of position velocity (PV) diagrams.
@@ 4.6.1 Stellar and gas kinematics around the W Nucleus
In this section we compare the gas and stellar kinematics around the W nucleus. We constructed a PV diagram of both the ionised gas and the hot molecular gas (Fig. 16), centered on the W nucleus with a PA 2576° directed along the major axis of rotation of the circumnuclear stellar disc calculated with the FIT KINEMATIC PA code (dashed black line in Fig. 15). We set the zero velocity at the redshift of the W nucleus (). The ionised gas (upper panel) reaches blueshifted velocities higher than those of the hot molecular gas (bottom panel). The stellar component (black star symbols in the plots) follows the kinematics of ionised gas, while it appears to rotate faster than the hot molecular gas.
In the PV diagram along the circumnuclear disc’s minor axis (PA 3476°), the ionised outflow (WNO), discussed in Sect. 4.4, is evident. It extends about 0.3” from the W nucleus and exhibits velocities reaching up to 600 km s-1 both in the blue- and red-shifted sides. However, this feature is not detected in the hot molecular gas, probably due to the dissociation of the H2 molecules or, more in general, to different physical conditions (e.g. in terms of magnetic field and electron density in the outflowing gas; see e.g. Kristensen et al. 2023).
We also constructed a PV diagram along the axis of the HG outflow (PA = 13°, left panel of Fig. 16), which is not exactly perpendicular to the axis of rotation of the circumnuclear disk. Ionised gas dominates the emission within 0.5” radius of the W nucleus, and decreases considerably at larger distances; instead, the H2 emission is more diffuse. The high blue- and red-shifted velocities spans a range of km s-1 km s-1 at a distance of about 1.5” (600 pc), delimiting the HG outflow.
To summarise, around the W nucleus we found 1) a relatively good match between stellar and gas kinematics, 2) a strong nuclear ionised outflow (blue and red-shifted) in addition to the HG outflow.
@@ 4.6.2 Stellar and gas kinematics around the E nucleus
Figure 17 shows the PV diagram along the major axis of the stellar rotational disc centered at the position of E nucleus (PA = 45°). Since it was not possible to decouple the contribution of the small-scale disc surrounding the E nucleus from the large-scale disc, the major axis of the small-scale disc was assumed to be aligned with the one of the large-scale disc computed in the previous subsection. We set the zero velocity at the redshift of the E nucleus ().
@@
A blue-shifted nuclear outflow is observed in both the hot molecular gas (bottom) and the ionized gas (top), reaching velocities of about km s-1. This emission forms the vertex of the SE conical outflow we isolated in Sect. 4.1.
@@ 5.1 Attenuation
Figure 18 shows the dust attenuation maps estimated from the different emission line ratios with a cut in S/N (¿3) for each line, as A. The attenuation AV varies between 2 and 14 mag, reaching higher values in the E nucleus (consistent with Perna et al. 2024) and lower values in the outer part of the FoV. The three attenuation maps computed with different line ratios are consistent. The attenuation computed from Eq. 1 integrating the fluxes in the central region (where the S/N of the faintest line Br ¿ 3) varies less than 20% between the three attenuation diagnostics. These values are also consistent with the stellar attenuation computed by Engel et al. (2011) through the K band, A, while our values are a factor 2-3 higher than their estimates through Pa/Br.
@@ 5.2 Star Formation
where C(Pa) = 3.90 assuming an upper mass cutoff of the IMF of 100 M⊙ and a solar metallicity. We derived a total star formation rate of 8 M⊙ yr-1. The value of SFR found with recombination line is significantly lower than the star formation computed with FIR and radio ( 200 M⊙ yr-1), possibly indicating that a large fraction of the SF is totally blocked by dust at NIR wavelengths (Varenius et al. 2016; Pereira-Santaella et al. 2021; Giménez-Arteaga et al. 2022; Perna et al. 2024).
The largest contribution to the emission of recombination lines comes from the elongated region located between the two nuclei with relatively faint H2 emission. In Fig. 19a we show the ratio of the total Pa flux to the total H2 flux, which peaks up to 30 north of the E nucleus and drops to values less than 1 in the outermost regions. We also see blobs in which the ratio reaches values of 2 (in log: 0.3-0.4 from Fig.19a); one of the largest is located about 1” west of the W nucleus. These blobs lie along a preferential NS direction, probably along a filament or dust lane where star formation is more intense.
Figure 19b shows the SFR surface density map, . The is highest in the elongated region between the two nuclei (10–40 M⊙ yr-1 kpc-2), while the other regions have value ¡ 1 M⊙ yr-1 kpc-2. The calculated over the entire FoV ( 1.5 kpc 1.5 kpc) has a value of 3 M⊙ yr-1 kpc-2. Based on FIR measurements, the SFR surface densities of the nuclear region of Arp 220 reaches values higher than 103 M⊙ yr-1 kpc-2. As discussed in Perna et al. (2024), this discrepancy may be explained taking into account different factors: the high obscuration caused by the dust, the different scale times of SF that the two bands (FIR and NIR) trace and to the possible presence of AGN that is completely obscured in the NIR. Typical star formation surface density values computed in radio band in ULIRGs are M⊙ yr-1 kpc-2 in the nuclear region ( kpc2), and M⊙ yr-1 kpc-2 on scales of few kpc (Lucatelli et al., 2024).
@@
Notes: The AGN luminosity refers to the nucleus from which the outflow originates. The SEO and NWB are attributed to the E nucleus, while the HGO and WNO to the W nucleus.

### v2

@@ 4.2 Hourglass outflow from the W nucleus
As shown in Fig. 1, both the ionized and molecular gas exhibit a shell-shaped morphology at high redshifted ( km s-1) velocities in the region north of the W nucleus. Additionally, the hot molecular gas shows a blueshifted shell-like structure in the region southwest of the same nucleus. The blue- and redshifted gas forms an ‘hourglass’ (HG)-like shape, compatible with an outflow launched from the W nucleus. This structure is even clearer in Fig. 5c, showing the sum of H2 spectral channel maps at km s-1 and km s-1 (with respect to the systemic velocity of the W nucleus), used to define our mask (white contours in the figure).
Within the HG region, we fitted simultaneously the hot molecular and the ionized gas using three Gaussian components, taking into account the large- and small-scale gas rotations as well as the outflow. To assign a Gaussian component to the HG outflow (HGO), we discarded the components with absolute velocities km s-1. Among the remaining ones, we attributed to the outflow the component with highest absolute velocity. The threshold (180 km s-1) represents a compromise: with a smaller velocity threshold, we would have included in the outflow also part of the rotational disk at larger scales; with a higher value, we would have missed part of the contribution of the outflow.
Figure 8 shows the flux, velocity and velocity dispersion of the gas component attributed to the outflow. The flux is dominated by the hot molecular gas over the ionized one. The velocity field is rather symmetrical with respect to the W nucleus and varies between and 400 km s-1. The highest observed velocity is located to the east in the redshifted shell, corresponding to an increase in molecular gas emission line flux. It is unclear whether this high-velocity gas is still part of the HGO, or whether it may represent a redshifted counterpart of the conical SE outflow associated with the E nucleus (Sect. 4.1).
@@
4.3 Northwest bubble
We identified a large scale bubble in the north west of the NIRSpec FoV, mostly visible in the [Fe ii] transitions (Fig. 1). Figure 5d displays the sum of spectral channel maps at km s-1 (with respect to the systemic velocity of E nucleus) of Pa and H2. We created the mask for the NW bubble (NWB) by selecting those spaxels with enhanced flux in these velocity channels, excluding the region around the W nucleus (white curve in the figure). This nuclear region was already included in Sect.4.2, but since its outflow kinematics are slightly different from those of the HG outflow, this region is discussed in Sect. 4.4.
The emission that contributes to the line profiles in this mask comes from the bubble and the large-scale rotating disk. Consequently, we fitted the emission lines with two Gaussian components. No specific boundaries on velocity and velocity dispersion were used during the fit. We assigned the most blueshifted component to the outflow. This Gaussian component also exhibits a higher velocity dispersion with respect to the one associated with the disk.
@@ 4.3 Northwest bubble
Figure 11 shows the fit of Pa and H2 in three spaxels extracted from the bubble. The black dashed curve represents the spectra from the original cube, highlighting the component (red-shaded area) attributed to the HGO that we have already isolated and subtracted. The red curves represent the NWB component, while the orange curves correspond to the other components used in the fit. In all panels, the nature of the Gaussian components is clear: the outflow component is blueshifted and broader than the disk component.
@@ 4.4 Compact outflow from the W nucleus
There is another kinematic feature in the W nuclear region that has not been isolated yet. It is a nuclear outflow (WNO) that is visible in the ionized gas phase (see also Perna et al. 2024). However, it affects just a few spaxels and cannot be isolated in the moment maps (Fig. 4) nor in the velocity channel maps (e.g., Fig. 1). Therefore, once all the features described before (i.e., the SEO, the HGO, and the NWB) have been subtracted, we fitted the residual cube with two components to model the WNO and the perturbed rotating gas (i.e., large-scale rotation over the entire FoV, or small-scale rotating nuclear discs).
@@
We attributed to the WNO the components with ¿ 150 km s-1 within a radius of 0.4” (150 pc) around the nucleus. In the more external regions, we associated both the components to the perturbed rotating gas. Figure 12 shows the three moments of the components attributed to the perturbed disk in H2 and Pa (top and central panels) and of the WNO (bottom panels). The WNO kinematics are characterized by blueshifted velocities up to km s-1 and a velocity dispersion of km s-1.
@@
Figure 13 shows the spectrum extracted from a random spaxel (marked in the bottom panel of Fig. 12) to illustrate the fit decomposition for the ionized and hot molecular gas phases. The differences in the line profile of the two gas phases are evident: the broad, blueshifted wing that is observed in Pa is not observed in the molecular gas.
4.5 Perturbed rotating disk
@@
The moment maps of hot molecular and ionized gas shown in Fig. 4 already suggested a large-scale rotation pattern, although strongly affected by several outflow features. These additional kinematic components were isolated and subtracted in Sects. 4.1, 4.2, 4.3 and 4.4. The fitting procedure used to separate the compact nuclear outflow (WNO) component from the rotating disk is described in the previous section (Sect. 4.4). After the subtraction of the components associated with the WNO, we assigned to the perturbed rotating gas all the components having km s-1. In fact, the components with higher velocity dispersion might represent residuals of outflowing gas that we did not model in the previous steps. Figure 12 (top and central panels) shows the moment maps of the isolated perturbed rotating gas, with contributions from both circum-nuclear discs in the vicinity of the E and W nuclei and the large-scale disk. The Moment-1 velocities are defined with respect to , which is more appropriate for the kiloparsec-scale disk (see Sect. 3.3).
@@
We observe a symmetric velocity gradient with respect to the E nucleus, with velocities from km s-1 (northeast) to km s-1 (southwest). The Moment-1 map in the northwest part of the FoV is instead disturbed due to the non-perfect subtraction of HGO and NWB contributions. Moreover, significant deviations from a large-scale rotational patter are also observed in the vicinity of the W nucleus: this is due to the presence of a small-scale disk surrounding the W nucleus (see also e.g., Scoville et al. 2017). We calculated the position angle (PA)111The position angle is measured on the receding half of the galaxy, taken anticlockwise from the north direction on the sky (east of north). of the major kinematic axis of the large-scale disk by fitting the Moment-1 map of Pa and for H2 with the FIT KINEMATIC PA code (Cappellari et al., 2007; Krajnović et al., 2011), obtaining 517° and 397°, respectively. In Sect. 4.6, we use the average of these two values to construct the position-velocity (PV) diagram.
@@
The Moment-2 of the perturbed rotating disk reaches values up to 150 km s-1, significantly lower than in the moment 2 map obtained before the subtraction of the outflowing features (Fig. 4). Moreover, the kinematics of Pa are more perturbed compared to that of the hot molecular gas, as also seen in Armus et al. (2023).
@@
Figure 14 shows three single-spaxel spectra extracted from the HG region (B), the SEO cone (C), and an overlapping region of the HG and NW bubble (D) (See Fig. 12). After the subtraction of the outflow components, the line profiles are now rather symmetrical; the second component used in the fit procedure (red dashed curves) either is much fainter than the first component, or shows similar velocity shifts and line widths. Therefore, the attribution of both components to the disk is well justified.
@@ 4.6 Stellar kinematics
Figure 15 presents the stellar kinematics maps obtained with pPXF. In particular, they are created from the analysis of the G235H cube, which is less affected by dust attenuation than the G140H cube and shows more prominent CO stellar absorption features in the wavelength range 2.2–2.4 m (Fig. 2). In Fig. 21 we show that the maps derived from the G235H and G140H gratings are consistent. As shown in the inset panel of Fig. 15, the stellar velocity field of Arp 220 presents two nuclear discs rotating in opposite directions separated by 1”, with a velocity offset of 120 km s-1 (see also Wheeler et al. 2020; Sakamoto et al. 1999). The stellar velocity dispersion is higher in the nuclear positions, and then decreases throughout a region crossing the two nuclei, in correspondence with the bright elongated region observed in the Pa Moment-0 map (Fig. 12). Moreover, in the E nucleus, the velocity dispersion increases toward the east, while in the W nucleus it increases in the south direction. In the outermost regions north of the two nuclei, the velocity dispersion reaches about 200 km s-1, consistent with the values found in Perna et al. (2020).
In the next two sub-sections we compare the stellar and gas kinematics making use of PV diagrams.
@@ 4.6.1 Stellar and gas kinematics around the W nucleus
In this section we compare the gas and stellar kinematics around the W nucleus. We constructed a PV diagram of both the ionized gas and the hot molecular gas (Fig. 16), centered on the W nucleus with a PA 2576° directed along the major axis of rotation of the circumnuclear stellar disk calculated with the FIT KINEMATIC PA code (dashed black line in Fig. 15). We set the zero velocity at the redshift of the W nucleus (). The ionized gas (upper panel) reaches blueshifted velocities higher than those of the hot molecular gas (bottom panel). The stellar component (black star symbols in the plots) follows the kinematics of ionized gas, while it appears to rotate faster than the hot molecular gas.
In the PV diagram along the circumnuclear disk’s minor axis (PA 3476°), the ionized outflow (WNO), discussed in Sect. 4.4, is evident. It extends about 0.3” from the W nucleus and exhibits velocities reaching up to 600 km s-1 both in the blue- and redshifted sides. However, this feature is not detected in the hot molecular gas, probably due to the dissociation of the H2 molecules or, more in general, to different physical conditions (e.g., in terms of magnetic field and electron density in the outflowing gas; see, e.g., Kristensen et al. 2023).
We also constructed a PV diagram along the axis of the HG outflow (PA = 13°, left panel of Fig. 16), which is not exactly perpendicular to the axis of rotation of the circumnuclear disk. Ionized gas dominates the emission within 0.5” radius of the W nucleus, and decreases considerably at larger distances; instead, the H2 emission is more diffuse. The high blue- and redshifted velocities spans a range of km s-1 km s-1 at a distance of about 1.5” (600 pc), delimiting the HG outflow. To summarize, around the W nucleus we found 1) a relatively good match between stellar and gas kinematics, 2) a strong nuclear ionized outflow (blue and redshifted) in addition to the HG outflow.
@@ 4.6.2 Stellar and gas kinematics around the E nucleus
Figure 17 shows the PV diagram along the major axis of the stellar rotational disk centered at the position of E nucleus (PA = 45°). Since it was not possible to decouple the contribution of the small-scale disk surrounding the E nucleus from the large-scale disk, the major axis of the small-scale disk was assumed to be aligned with the one of the large-scale disk computed in the previous subsection. We set the zero velocity at the redshift of the E nucleus ().
@@
A blueshifted nuclear outflow is observed in both the hot molecular gas (bottom) and the ionized gas (top), reaching velocities of about km s-1. This emission forms the vertex of the SE conical outflow we isolated in Sect. 4.1.
@@ 5.1 Attenuation
Figure 18 shows the dust attenuation maps estimated from the different emission line ratios with a cut in S/N (¿3) for each line, as A. The attenuation AV varies between 2 and 14 mag, reaching higher values in the E nucleus (consistent with Perna et al. 2024) and lower values in the outer part of the FoV. The three attenuation maps computed with different line ratios are consistent. The attenuation computed from Eq. 1 integrating the fluxes in the central region (where the S/N of the faintest line Br ¿ 3) varies less than 20% between the three attenuation diagnostics. These values are also consistent with the stellar attenuation computed by Engel et al. (2011b) through the K band, A, while our values are a factor 2-3 higher than their estimates through Pa/Br.
@@ 5.2 Star formation
where C(Pa) = 3.90 assuming an upper mass cutoff of the IMF of 100 M⊙ and a solar metallicity. We derived a total SFR of 8 M⊙ yr-1. The value of SFR found with the recombination line is significantly lower than ones computed in a similar area with FIR and radio data ( 200 M⊙ yr-1, from Varenius et al. 2016 and Pereira-Santaella et al. 2021), possibly indicating that a large fraction of the SF is totally blocked by dust at NIR wavelengths (Giménez-Arteaga et al. 2022; Perna et al. 2024).
The largest contribution to the emission of recombination lines comes from the elongated region located between the two nuclei with relatively faint H2 emission. In Fig. 20a we show the ratio of the total Pa flux to the total H2 flux, which peaks up to 30 north of the E nucleus and drops to values less than one in the outermost regions. We also observed blobs in which the ratio reaches values of approximately two (in log: 0.3-0.4 from Fig.20a); one of the largest is located about 1” west of the W nucleus. These blobs lie along a preferential NS direction, probably along a filament or dust lane where SF is more intense.
Figure 20b shows the SFR surface density map, . The is highest in the elongated region between the two nuclei (10–40 M⊙ yr-1 kpc-2), while the other regions have value ¡ 1 M⊙ yr-1 kpc-2. The calculated over the entire FoV ( 1.5 kpc 1.5 kpc) has a value of 3 M⊙ yr-1 kpc-2. Similar values are reported in a sample of ULIRGs studied with VLT/SINFONI (Piqueras López et al. 2012), ranging between 0.4 and 2 M⊙ yr-1 kpc-2. Based on FIR measurements, the SFR surface densities of the nuclear region of Arp 220 reaches values higher than 103 M⊙ yr-1 kpc-2. As discussed in Perna et al. (2024), this discrepancy may be explained taking into account different factors: the high obscuration caused by the dust, the different scale times of SF that the two bands (FIR and NIR) trace and to the possible presence of AGN that is completely obscured in the NIR. Typical star formation surface density values computed in radio band in ULIRGs are M⊙ yr-1 kpc-2 in the nuclear region ( kpc2), and M⊙ yr-1 kpc-2 on scales of few kiloparsec (Lucatelli et al., 2024).
@@
Notes: (1) and (2) SFR computed from radio/FIR by Varenius et al. (2016) and Pereira-Santaella et al. (2021), and from Pa by Perna et al. (2024), respectively. (3) AGN luminosity from IR. (4) AGN luminosity from X-rays by Paggi et al. (2017). (5) Hot-to-cold ratio. (6) Outflow radius. (7) [Fe ii]-based electron density. (8) Outflow velocity, calculated as the flux-weighted average of the velocity (defined as in Eq. 6) of each spaxel and the associated errors. (9) Dynamical time. (10) Average attenuation. (11) Outflow mass. (12) Outflow kinetic energy. (13) Outflow kinetic power. (14) Outflow mass loading factor.
The AGN luminosity refers to the nucleus from which the outflow originates. The SEO and NWB are attributed to the E nucleus, while the HGO and WNO to the W nucleus.